OpenStudy (caozeyuan):

A newly discovered planet has a mean radius of 2020 km. A vehicle on the planet\'s surface is moving in the same direction as the planet\'s rotation, and its speedometer reads 145 km/h. If the angular velocity of the vehicle about the planet\'s center is 5.28 times as large as the angular velocity of the planet, what is the period of the planet\'s rotation?

OpenStudy (caozeyuan):

@perl

OpenStudy (perl):

hey

OpenStudy (caozeyuan):

can you help me?

OpenStudy (perl):

yes

OpenStudy (perl):

lets call the period of the planet to make a complete revolution T

OpenStudy (perl):

lets write some stuff down

OpenStudy (perl):

the angular velocity of planet 2pi / Tp angular velocity of vehicle 2pi / Tv tangential velocity = radius * angular velocity

OpenStudy (surry99):

From v = omega x r you can get omega (angular velocity for the vehicle. Then, you can calculate omega for the planet because omega vehicle/omega planet = 5.28 Once you have omega for the planet, omega = 2 x pi x frequency Solve for the frequency and then f = 1/T where T is the period

OpenStudy (perl):

Tv stands for period of vehicle Tp stands for period of planet so we are given 2 Pi / (Tv) = 5 * 2 Pi / (Tp)

OpenStudy (perl):

and tangential velocity is 145 km /hr

OpenStudy (perl):

so we can find Tv 145 = (2020 ) * 2Pi / Tv once we find Tv, we can substitute to find Tp

OpenStudy (perl):

Tv = 87.53 hours is the period of the vehicle

OpenStudy (perl):

we were given 2 Pi / (Tv) = 5 * 2 Pi / (Tp) so plug in and solve for Tp

OpenStudy (perl):

with me so far?

OpenStudy (perl):

Tp = 437.66 hours

OpenStudy (caozeyuan):

the system says incorrect

OpenStudy (perl):

does it say to round

OpenStudy (caozeyuan):

no, only the word "incorrect" appears

OpenStudy (perl):

oh i multiplied by 5, not 5.28

OpenStudy (perl):

462.165

OpenStudy (caozeyuan):

no, still not right, I'll skip it for now

OpenStudy (perl):

i did it also the other way the person above me 462.165 hours is the period of the planet

OpenStudy (caozeyuan):

Incorrect. Note that the speedometer only reports the speed of the vehicle relative to the ground. now there is something more

OpenStudy (perl):

@surry99 please check

OpenStudy (perl):

145 = 2020 * w_v 145/2020 = w_v w_p = w_v/5.28 w_p = (145/2020) / 5.28 T = 2pi / (w_p)

OpenStudy (perl):

T = 462.1651118

OpenStudy (surry99):

Ah yes, forgot about what the speedometer measures ...silly me. hang on.

OpenStudy (surry99):

Is the answer 374 hours

OpenStudy (surry99):

V vehicle = V planet + V vehicle relative to planet r x omega v = r omega p + 145 and omega vehicle = 5.28 x omega planet so: r x (5.28 omega p) = r x omega p + 145 solving for omega p... omega p = 145/(4.28 x r) we get omega p = 145 km/hr/(4.28 x 2020 km) = .0167/hr then T = 2 x pi/omega p = 6.28/.0167 /hr = 376 hours

OpenStudy (perl):

hmm

OpenStudy (perl):

thankyou surry for your help @surry99

OpenStudy (surry99):

It was a pleasure working with you @perl